# Context Parametrization with Compositional Adapters

### v1

@@ 1 Introduction
Large language models (LLMs) adapt to new tasks by integrating contextual information pertaining to these tasks, such as instructions, demonstrations, or extracted evidence. This adaptability is typically realized through in-context learning (ICL) (Brown et al., 2020), instruction following (Ouyang et al., 2022), and retrieval-augmented generation (Lewis et al., 2020; Borgeaud et al., 2022), where extra tokens in the prompt act as a transient memory that steers model behavior. Another common strategy is to adapt parameters directly through supervised fine-tuning (SFT), ranging from full model updates (Devlin et al., 2019; Raffel et al., 2020) to parameter-efficient variants Pfeiffer et al. (2023). While effective, both prompting and SFT face limitations. Prompt-based methods often require long contexts, which destabilize attention and degrade performance as length increases (Liu et al., 2023; Kossen et al., 2023), while also incurring inference costs that scale with context size. SFT, on the other hand, requires additional training overhead and lacks flexibility across tasks.
A complementary line of work views adapters (Houlsby et al., 2019; Hu et al., 2022) as a mechanism for parametrizing context. In our work, we use context as an umbrella term for instructions, demonstrations, or supporting passages. By translating context into parameters, adapters offer a persistent representation that replaces prompt tokens, thereby improving the stability of processing long contexts and amortizing adaptation across tasks (Karimi Mahabadi et al., 2021; He et al., 2022; Liu et al., 2023). Once context is encoded into adapters, inference requires only the query tokens, reducing the cost of processing long prompts and making the approach especially attractive in latency- or memory-constrained settings. This context-to-adapter transformation encodes an otherwise transient token sequence into a manipulable object, allowing for caching, reuse, and algebraic combination.
Building on this perspective, recent work has leveraged meta-learning to efficiently generate adapter parameters directly from support examples without fine-tuning anew for each context (Zhang et al., 2023; Wang et al., 2023; Chen et al., 2025).
@@
To bridge this critical gap, we introduce CompAs (Compostional Adapters), a framework that maps context into adapter parameters with compositional structure at its core. See Figure 1 for a high-level overview. In CompAs, a teacher LM process concatenated contexts, while the student LM learns to approximate this behavior using only query tokens, complemented with the sum of adapters generated for individual contexts. This is facilitated by a generator network, which maps individual contexts into compositional adapters. CompAs is driven by auxiliary compositionality and reconstruction losses, which provide diagnostics for whether parameter-space operations preserve semantics and faithfulness (Jacovi & Goldberg, 2020; Turpin et al., 2023; Min et al., 2023; Lampinen et al., 2022). We evaluate CompAs on multiple-choice and extractive QA tasks, where either demonstrations (MMLU, ARC-Challenge) or retrieved passages (HotpotQA) serve as context. We observe consistent improvements over ICL, fine-tuning and existing context-to-parameter methods. These gains are most pronounced when scaling to larger numbers of demonstrations, highlighting the capacity of CompAs for efficient integration of external evidence by composition in parameter-space.
@@ 4.1 Experimental Setup
We experiment with decoder-only LMs: LLaMA-3.1 8B (LLaMA 8B) and Llama-3.2 3B (Dubey et al., 2024), and Qwen-2.5 7B (Yang et al., 2025). For brevity, we refer to these models as LLaMA 8B, LLaMA 3B, and Qwen 7B.
@@ Methods.
We compare against three representative approaches. ICL is standard $k$-shot prompting with demonstrations concatenated in prompt text. Generative Adapter (GenAda) (Chen et al., 2025) employs a nonlinear hypernetwork to predict adapter weights from context, trained with reconstruction and continuation losses. In contrast, our method enforces a linear bottleneck, generating adapters via a lightweight adapter module followed by a linear transformation. wilda (Jukić & Šnajder, 2025) fine-tunes a separate adapter for each context, achieving strong accuracy but incurring heavy computational overhead, since a new adapter must be trained for every context.
@@ 6 Related Work
Analyses show ICL often exploits surface-level cues rather than learning task semantics (Min et al., 2022), and that transformers can internally implement simple learning algorithms during inference (Xie et al., 2021; Akyürek et al., 2022). In parallel, parameter-efficient fine-tuning (PEFT) replaces long prompts with compact trainable modules such as soft/prefix prompts (Li & Liang, 2021; Lester et al., 2021) or low-rank adapters (Hu et al., 2021). Recently, Hong et al. (2024) proposed mixtures of in-context learners to efficiently combine subsets of demonstrations. Like PEFT, our work keeps inference on the query only, but instead of concatenating demonstrations, we translate them into compositional adapters.
@@
Another line of research explores how to compose learned modules or entire models. In the adapter setting, AdapterFusion learns to fuse multiple task adapters non-destructively (Pfeiffer et al., 2021), while MAD-X composes language and task adapters for cross-lingual transfer (Pfeiffer et al., 2020). For full model weights, model soups average fine-tuned checkpoints to improve robustness (Wortsman et al., 2022), and task vectors use weight differences to add or subtract task behaviors (Ilharco et al., 2022). These approaches combine pre-trained modules or models after the fact. In contrast, we generate adapters on-the-fly from context and train them explicitly for additivity, so that parameter-space composition mirrors text concatenation.
@@
HyperNetworks generate parameters of one network using another (Ha et al., 2017). Ansell et al. (2021) introduced MAD-G, where a hypernetwork generates language adapters conditioned on language embeddings. Subsequent work extended this idea to few-shot adaptation (Zhang et al., 2023; Chen et al., 2025). Our method can be viewed as a hypernetwork explicitly designed for compositionality: a generator maps support encodings into LoRA parameters, while compositional distillation and a linearity regularizer enforce that adapter addition in parameter space approximates context concatenation in input space. This aligns with broader efforts to impose linear additive structure on representations (Sulc, 2025), but uniquely emphasizes parameter-space composition as a scalable route to context integration.
@@ B.1 Data
For MMLU (Hendrycks et al., 2021), which comprises $59$ subject-specific subsets, we report accuracy averaged across all subsets. For weak supervision during generator optimization, we sample queries and contexts from the validation split without using ground-truth labels, while demonstrations are drawn (with labels) from the training split. Evaluation is performed on the held-out test split.
@@ B.3 Methods
For GenAda (Chen et al., 2025) and wilda (Jukić & Šnajder, 2025), we use the default hyperparameters and follow the training procedures described in the respective papers. To ensure comparability, we adopt an identical configuration for adapter sizes and target modules across all methods; these design choices are detailed in the following section.
@@ Appendix C Additional Experiments
C.1 Positional Encoding and Order Sensitivity
@@
C.2 Efficiency
@@ Appendix E Limitations
In preparing this manuscript, we made limited use of LLMs to polish the writing at the sentence level. This assistance was limited to stylistic improvements such as grammar, clarity, and conciseness.

### v2

@@ 1 Introduction
Large language models (LLMs) adapt to new tasks by integrating contextual information pertaining to these tasks, such as instructions, demonstrations, or extracted evidence. This adaptability is typically realized through in-context learning (ICL) (Brown et al., 2020), instruction following (Ouyang et al., 2022), and retrieval-augmented generation (Lewis et al., 2020; Borgeaud et al., 2022), where extra tokens in the prompt act as a transient memory that steers model behavior. Another common strategy is to adapt parameters directly through supervised fine-tuning (SFT), ranging from full model updates (Devlin et al., 2019; Raffel et al., 2020) to parameter-efficient variants Pfeiffer et al. (2023). While effective, both prompting and SFT face limitations. Prompt-based methods often require long contexts, which destabilize attention and degrade performance as length increases (Liu et al., 2024), while also incurring inference costs that scale with context size. SFT, on the other hand, requires additional training overhead and lacks flexibility across tasks.
A complementary line of work views adapters (Houlsby et al., 2019; Hu et al., 2022) as a mechanism for parametrizing context. In our work, we use context as an umbrella term for instructions, demonstrations, or supporting passages. By translating context into parameters, adapters offer a persistent representation that replaces prompt tokens, thereby improving the stability of processing long contexts and amortizing adaptation across tasks (Mahabadi et al., 2021; He et al., 2022; Liu et al., 2024). Once context is encoded into adapters, inference requires only the query tokens, reducing the cost of processing long prompts and making the approach especially attractive in latency- or memory-constrained settings. This context-to-adapter transformation encodes an otherwise transient token sequence into a manipulable object, allowing for caching, reuse, and algebraic combination.
Building on this perspective, recent work has leveraged meta-learning to efficiently generate adapter parameters directly from support examples without fine-tuning anew for each context (Bansal et al., 2022; Chen et al., 2025).
@@
To bridge this critical gap, we introduce CompAs (Compostional Adapters), a framework that maps context into adapter parameters with compositional structure at its core. See Figure 1 for a high-level overview. In CompAs, a teacher LM process concatenated contexts, while the student LM learns to approximate this behavior using only query tokens, complemented with the sum of adapters generated for individual contexts. This is facilitated by a generator network, which maps individual contexts into compositional adapters. CompAs is driven by auxiliary compositionality and reconstruction losses, which provide diagnostics for whether parameter-space operations preserve semantics and faithfulness (Jacovi and Goldberg, 2020; Turpin et al., 2023; Lampinen et al., 2022). We evaluate CompAs on multiple-choice and extractive QA tasks, where either demonstrations (MMLU, ARC-Challenge) or retrieved passages (HotpotQA) serve as context. We observe consistent improvements over ICL, fine-tuning and existing context-to-parameter methods. These gains are most pronounced when scaling to larger numbers of demonstrations, highlighting the capacity of CompAs for efficient integration of external evidence by composition in parameter-space.
@@ 4.1 Experimental Setup
We experiment with decoder-only LMs: LLaMA-3.1 8B (LLaMA 8B) and Llama-3.2 3B (Grattafiori et al., 2024), and Qwen-2.5 7B (Yang et al., 2025). For brevity, we refer to these models as LLaMA 8B, LLaMA 3B, and Qwen 7B.
@@ Methods.
We compare against three representative approaches. ICL is standard $k$-shot prompting with demonstrations concatenated in prompt text. Generative Adapter (GenAda) (Chen et al., 2025) employs a nonlinear hypernetwork to predict adapter weights from context, trained with reconstruction and continuation losses. In contrast, our method enforces a linear bottleneck, generating adapters via a lightweight adapter module followed by a linear transformation. wilda (Jukić and Šnajder, 2025) fine-tunes a separate adapter for each context, achieving strong accuracy but incurring heavy computational overhead, since a new adapter must be trained for every context.
@@ 6 Related Work
Analyses show ICL often exploits surface-level cues rather than learning task semantics (Min et al., 2022), and that transformers can internally implement simple learning algorithms during inference (Xie et al., 2022; Akyürek et al., 2023). In parallel, parameter-efficient fine-tuning (PEFT) replaces long prompts with compact trainable modules such as soft/prefix prompts (Li and Liang, 2021; Lester et al., 2021) or low-rank adapters (Hu et al., 2022). Recently, Hong et al. (2025) proposed mixtures of in-context learners to efficiently combine subsets of demonstrations. Like PEFT, our work keeps inference on the query only, but instead of concatenating demonstrations, we translate them into compositional adapters.
@@
Another line of research explores how to compose learned modules or entire models. In the adapter setting, AdapterFusion learns to fuse multiple task adapters non-destructively (Pfeiffer et al., 2021), while MAD-X composes language and task adapters for cross-lingual transfer (Pfeiffer et al., 2020). For full model weights, model soups average fine-tuned checkpoints to improve robustness (Wortsman et al., 2022), and task vectors use weight differences to add or subtract task behaviors (Ilharco et al., 2023). These approaches combine pre-trained modules or models after the fact. In contrast, we generate adapters on-the-fly from context and train them explicitly for additivity, so that parameter-space composition mirrors text concatenation.
@@
HyperNetworks generate parameters of one network using another (Ha et al., 2017). Ansell et al. (2021) introduced MAD-G, where a hypernetwork generates language adapters conditioned on language embeddings. Subsequent work extended this idea to few-shot adaptation (Chen et al., 2025). Our method can be viewed as a hypernetwork explicitly designed for compositionality: a generator maps support encodings into LoRA parameters, while compositional distillation and a linearity regularizer enforce that adapter addition in parameter space approximates context concatenation in input space. This aligns with broader efforts to impose linear additive structure on representations, but uniquely emphasizes parameter-space composition as a scalable route to context integration.
@@ B.1 Data
For MMLU (Hendrycks et al., 2021), which comprises $59$ subject-specific subsets, we report accuracy averaged across all subsets. We train a single generator jointly across all subsets. This ensures that the generator is shared across domains rather than tuned separately per subject. For weak supervision during generator optimization, we sample queries and contexts from the validation split without using ground-truth labels, while demonstrations are drawn (with labels) from the training split. Evaluation is performed on the held-out test split.
@@ B.3 Methods
For GenAda (Chen et al., 2025) and wilda (Jukić and Šnajder, 2025), we use the default hyperparameters and follow the training procedures described in the respective papers. To ensure comparability, we adopt an identical configuration for adapter sizes and target modules across all methods; these design choices are detailed in the following section.
@@ Appendix C Additional Experiments
C.1 Long-Form Summarization
Beyond question answering, we also evaluate whether the representations learned by CompAs generalize to long-form generation tasks. Its reconstruction objective already requires producing full demonstrations and multi-paragraph passages (Table 3), exercising long-range generation in training.
To test generalization capabilities beyond QA, we additionally evaluate CompAs on XSUM, a standard abstractive summarization benchmark. Results are reported in Table 6. CompAs achieves the strongest performance across all ROUGE metrics, substantially outperforming ICL and other baselines. This indicates that the learned context-to-parameter mapping transfers to tasks that require extended, structured generation, supporting the use of CompAs as a general parametrization mechanism.
Table 6: XSUM summarization with LLaMA-3.1 8B. We report ROUGE-1 (R1), ROUGE-2 (R2), and ROUGE-L (RL) F1. R1 and R2 measure unigram and bigram overlap with the reference summary, while RL measures the longest common subsequence. Higher values indicate closer alignment with the reference and better summary quality.
Method R1 $\uparrow$ R2 $\uparrow$ RL $\uparrow$
ICL 36.1 15.6 28.4
PBFT 38.3 18.1 31.6
GenAda 33.4 13.1 25.6
CompAs 41.8 22.8 33.5
C.2 Supervised and Prompt-Based Baselines
We further evaluate supervised and parametric baselines: (i) pattern-based fine-tuning (PBFT), optimized with a cloze-style language modeling objective (Schick and Schütze, 2021); (ii) a linear SFT head trained on top of the frozen backbone; (iii) prefix-tuning with the same demonstrations (Li and Liang, 2021); and (iv) an ICL variant that retrieves the top-$k$ most similar demonstrations for each query using SBERT embeddings (Reimers and Gurevych, 2019).
Results for MMLU and ARC are reported in Table 7. Across both benchmarks, CompAs consistently outperforms all supervised and prompt-based baselines, supporting that its improvements arise from the context-to-parameter mechanism rather than differences in data, supervision, or training procedure.
Table 7: Supervised and prompt-based baselines with LLaMA-3.1 8B. We report mean accuracy on MMLU and ARC over $10$ runs with $16$ demonstrations in total. CompAs is evaluated with $4\times 4$ composition. Standard deviations are shown in subscript, and the best results are in bold.
Method MMLU ARC
ICL (16-shot) $64.2_{1.8}$ $74.2_{1.6}$
ICL + prefix-tuning $66.1_{1.4}$ $74.0_{1.3}$
ICL + similarity selection $68.1_{0.6}$ $75.5_{0.2}$
PBFT (labels only) $66.8_{1.1}$ $75.6_{1.0}$
Linear SFT head $62.1_{1.5}$ $71.3_{1.4}$
CompAs ($\times4$) $\mathbf{72.2_{0.3}}$ $\mathbf{81.3_{0.3}}$
C.3 Positional Encoding and Order Sensitivity
@@
C.4 Compositional Scaling
A central design choice in CompAs is how context is partitioned into blocks prior to adapter generation. Block size controls both how much information is encoded into a single adapter and how many such adapters must be composed at inference time, and therefore directly influences both expressiveness and compositional stability. We next examine how this tradeoff affects accuracy and additivity as the total number of demonstrations increases.
Table 9 shows how block structure affects both task accuracy and compositional fidelity. For a fixed number of demonstrations, CompAs reaches its best performance with intermediate block sizes (approximately 4–8 demonstrations per block), which yield both higher MMLU accuracy and lower additivity error $\eta$.
When blocks become too large (e.g., $2\times 16$ or $1\times 16$), the generator must encode many demonstrations jointly, which weakens the linear structure of the latent parameterization and increases additivity error. Conversely, when blocks become too small (e.g., $16\times 1$), each block provides a limited contextual signal, leading to slightly reduced accuracy and moderately higher $\eta$.
As the total number of demonstrations increases, accuracy steadily improves, demonstrating that CompAs benefits from additional evidence. In practice, the maximum feasible block size is limited by GPU memory, since larger blocks require the generator to encode more tokens jointly. Within these limits, using moderately sized blocks provides a favorable tradeoff between representation stability and information density.
Finally, long-context ICL baselines can only be evaluated reliably up to $32$ demonstrations due to memory constraints. In contrast, CompAs decouples inference cost from context length, enabling evaluation with more demonstrations on the same hardware through compositional adapter merging.
Table 9: Effect of block granularity and scaling on MMLU accuracy and additivity error $\eta$ with LLaMA 8B. We vary the block structure for a fixed total number of demonstrations and report accuracy and generator additivity error.
Total demos Method Accuracy $\bm{\eta}$ $\downarrow$
16 ICL (16-shot) 64.2
16 CompAs ($\times4$) 72.2 0.08
[…]
16 CompAs ($\times1$6) 69.4 0.15
[…]
32 ICL (32-shot) 65.5
[…]
64 CompAs ($\times8$) 76.2 0.16
C.5 Robustness to Noisy Context
In practical retrieval-based systems, the contexts provided to a model are often imperfect: demonstrations may contain incorrect labels, multiple sources may disagree about the same example, or retrieved passages may be unrelated to the query. We therefore evaluate how CompAs behaves under controlled perturbations that explicitly simulate these failure modes and compare it directly to standard ICL under identical corruption patterns.
We introduce three types of degradation at inference time. Noisy demonstrations are generated by randomly flipping the labels of a fixed fraction of the in-context examples (12.512.5% and 2525% of the $16$ demonstrations), while keeping inputs unchanged. This simulates annotation errors or unreliable retrieval from weakly supervised corpora. Contradictory demonstrations are constructed by duplicating a subset of examples and assigning them conflicting labels: for $4$ of the $16$ demonstrations, the same input appears twice across contexts, once with the correct label and once with an incorrect label. This explicitly creates inconsistent supervision inside the context and mimics disagreement between retrieved sources or conflicting knowledge bases. Off-topic contamination is introduced by appending unrelated paragraphs sampled from HotpotQA to the end of the context, increasing the number of irrelevant tokens without changing the gold demonstrations. This simulates retrieval failures in which non-informative or spurious documents are mixed into the prompt. All results are reported as averages over 10 independent runs with different random corruption patterns.
Table 10 reports accuracy on MMLU as the demonstration noise increases. Both methods degrade as noise grows, but CompAs consistently retains a clear performance margin. Notably, the gap widens under stronger label corruption and remains substantial under contradictory supervision. This suggests that compositional context encoding is more robust to internally inconsistent evidence than raw concatenation: while ICL treats all tokens uniformly, CompAs compresses each demonstration into a latent representation, reducing the impact of spurious or conflicting signals.
Table 11 reports results on HotpotQA as irrelevant passages are injected into the context. ICL performance degrades steadily as more off-topic content is introduced, indicating sensitivity to prompt dilution and distraction effects. In contrast, CompAs remains comparatively stable and preserves most of its clean-context accuracy even as irrelevant text grows. This behavior is consistent with the hypothesis that encoding context into parameters rather than raw tokens makes inference less sensitive to unrelated input.
Taken together, these experiments demonstrate that composition-based context parametrization improves robustness to realistic retrieval noise. Even when demonstrations contain errors, contradictions, or irrelevant information, CompAs degrades more gracefully than ICL, supporting additive composition as a stabilizing mechanism under imperfect and noisy context conditions.
Table 10: Robustness under noisy and contradictory demonstrations on MMLU with LLaMA 8B (mean over $10$ runs).
Method Clean 12.5% noise 25% noise Contradictory
ICL 66.5 62.1 58.7 57.9
CompAs 72.2 71.3 69.4 68.4
Table 11: HotpotQA under off-topic context contamination with LLaMA 8B. Accuracy with irrelevant paragraphs appended to the context (mean over $10$ runs).
Method Clean +1 off-topic +2 off-topic
ICL 82.0 79.4 77.6
CompAs 82.9 82.3 81.7
C.6 Efficiency
@@
Additionally, Table 13 compares absolute inference and training compute across methods. The methods differ primarily in their training regimes. WILDA is inexpensive for a single demonstration configuration, but does not amortize: each new context requires a separate fine-tuning run. GenAda incurs a large one-time training cost to learn a generator. CompAs lies between these extremes, training a generator once at moderate cost and reusing it across contexts. In cumulative terms, WILDA’s total cost grows linearly with the number of demonstration configurations and overtakes CompAs after approximately $17$ contexts, beyond which repeated fine-tuning becomes more expensive than the one-time generator training of CompAs. At the same time, CompAs is approximately $13\times$ cheaper to train than GenAda, while achieving comparable inference-time efficiency. Thus, CompAs provides a favorable tradeoff between repeated fine-tuning and large upfront training cost, combining amortization with moderate training overhead.
@@
Table 13: Absolute compute comparison with LLaMA 8B. We report absolute inference-time compute per query and training compute in TFLOPs on the ARC-Challenge dataset using $16$ demonstrations in total with $4\times 4$ composition for CompAs. WILDA trains a new adapter for each demonstration configuration, whereas GenAda and CompAs train a generator once and amortize across different contexts.
Method Inference TFLOPs / query Training TFLOPs Regime ICL (prompting) $4.0$ No training WILDA $0.7$ $1.4\times 10^{5}$ Per demo set GenAda $0.9$ $3.1\times 10^{7}$ One-time CompAs $0.8$ $2.4\times 10^{6}$ One-time
@@ Appendix E Limitations
In preparing this manuscript, we used LLMs to polish the writing at the sentence level. This assistance was limited to stylistic improvements such as grammar, clarity, and conciseness.